OpenStudy (anonymous):

determine the slope of the line passing through each pair of points: (8,-4) and (6,1)

OpenStudy (katrinakaif):

Fidn teh slope of these points by using the equation: (y2-y1)/(x2-x1) Then plug the slope in as "m" in the following equation: (y-y1) = m(x-x1) Plug in m and then Pick one of the points to plug in as (x1,y1) Solve =)

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(1-(-4))/(6-8) m=(1+4)/(6-8) m=(5)/(6-8) m=(5)/(-2) m=-5/2

OpenStudy (anonymous):

is that a fraction? or division?

jimthompson5910 (jim_thompson5910):

it's the fraction \[\large -\frac{5}{2}\]